# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

### v1

@@ Abstract
The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time that started in August 2023 and lasted for almost 6 months. We study the timing and spectral properties of this source using publicly available archival Insight-HXMT data for the first 10 observation IDs that last from MJD 60181 to 60198 with a total of 92 exposures for all three energy bands. We extracted the quasi-periodic oscillation properties by model fitting the power density spectrum and from those properties we designate that the QPOs are type-C in nature. We also conclude that the origin of the QPOs could be the shock instabilities, formed in the transonic accretion flow around black holes. The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of HXMT in the broad energy band of $2-150$ keV. To achieve the best fit, we needed a combination of interstellar absorption, power-law, multi-color disk-blackbody continuum, gaussian emission/absorption, and power-law reflection by neutral material. From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was making a transition toward the softer states. The shock (the boundary layer of the corona) moved inward in progressive days in accordance with the spectral nature. We found that the source is present in a high-inclination binary system. The hydrogen column density was found with an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
@@ 1 Introduction
Black holes (BHs) are one of the end-products of massive stars. Stellar-mass black holes (SBHs) are one of the three types of presently detected BHs. They reside in binary systems with a companion star and most of their radiation comes out in the form of X-rays. Thus, they are called black hole X-ray binaries or BHXRBs. Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs). LMXRBs consist of type-A or later type of stars whereas the HMXRBs consist of giant O or B-type of stars (White et al. 1995; Tetarenko et al. 2016). Matter coming from the companion accumulates at a distance from the BH due to lack of viscosity, known as the pile-up radius (Chakrabarti et al. 2019; Chatterjee et al. 2022). Due to the accumulation of more matter, the temperature will rise. This gradual increment of temperature sets in instability at the pile-up, which therefore increases the viscosity. At some point in time when the accumulated matter will gain enough viscosity to push forward, accretion starts. On the other hand, irradiation of the accumulated matter at the outer boundary by the central object can also trigger mass accretion when the temperature of the accumulated matter crosses a critical limit (King & Ritter 1998; Mondal 2020). Thus, these systems make ideal laboratories for studying the physics of accretion. The BHXRBs mostly stay in a quiet dormant state when their flux stays at a very low detection level. Episodically, the radiation level increases by a very large factor and they become easily detectable. This is known as an outburst (Remillard & McClintock 2006). The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence. Depending on the nature of the outburst, the sources are classified into two types: persistent and transient sources. Most of the BHXRBs are transient type of sources. They stay in the dormant state for most of their lifetimes. The luminosity becomes $\sim 10^{35}$ erg s-1 during episodic outbursting phase.
@@
During the onset of an outburst, the flux in the light curve changes noticeably. According to nature, the outbursts are divided into two types (Debnath et al. 2010): fast rise slow decay (FRSD) and slow rise slow decay (SRSD). However, there are also some other classified outburst types, e.g., according to Zhang et al. (2019), outbursts can also be classified into glitch, reflare, multipeak, and mini-outburst. There are generally four designated spectral states during a complete BH outburst (Remillard & McClintock 2006). They are known as hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS). Generally, when an outburst starts, it starts in the HS. Then it slowly moves toward the SS through the HIMS and SIMS (Belloni et al. 2005, 2011, and references within). After it reaches its SS the rising phase generally comes to an end and the outburst starts its decaying phase. In the decaying phase, it goes in the opposite direction and reaches the HS again. When a BH outburst goes through all these states, it is known as a type-I outburst. When the soft state is absent, it is known as a ‘failed’ or type-II outburst.
The radiation spectrum of a BH generally consists of two main components. One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973). The other is the hard power-law component, which extends up to very high energies. The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which are intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al. 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al. 1999). When the outburst is in HS, there is very small contribution from the thermal component and mostly the non-thermal component dominates. In SS, the thermal component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable. other than these two components, other components could also be present in the radiation spectrum of a BH. In the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a very close distance to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Psaltis & Lamb 1997; Borozdin et al. 1999; Psaltis 2001;). It was been recently suggested by Chatterjee et al. (2023) that BMC was present in the radiation spectrum of the BHC MAXI J0637-430. The reprocessed emission plays an important role in the case of BHXRBs. Fraction of electromagnetic radiation can come back from the optically thin and geometrically thick Compton cloud to the optically thick and geometrically thin accretion disk due to the strong gravitational bending effect of the BH (Riaz et al. 2021). The illumination of the Comptonized photons from the Compton cloud to the accretion disk is known as returning or reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010). The reflection spectrum is a very important part of the whole radiation spectrum which helps determine many important properties of the source. The reflection radiation can produce some very specific features in the spectrum. One of which is the iron $K_{\alpha}$ emission line at $\sim 6.4$ keV. This happens when the returning radiation illuminates the disk and there is the fluorescence of ionized Fe. This line emission has a width of the order of $\sim 1~{}eV$ (Pozdniakov et al. 1979). However, for rotating black holes there are effects like Doppler shift and gravitational broadening which makes the line broadened and for rapidly rotating black holes this line can have two peaks due to red-shifting and blue-shifting of the component. This line emission is a very important aspect in estimating the spin of the BH. There could also be other features like Fe-K absorption edge and Compton hump. These features arise because of counterplay between two phenomena. Which feature will be prominent will be resulted in depending on which phenomenon is dominant. These are also energy-dependent. The interplay between photoelectric absorption and Compton scattering phenomena (Guilbert & Rees 1988; Lightman & White 1988; Fabian et al. 1989) will decide the spectral feature. At higher energies above $20$ keV, the scattering will be dominant resulting in a Compton hump, whereas, in low energies, there will be the dominance of photoelectric absorption and the spectrum drops away (Lightman & White 1988). As these features take place at a closer distance from the compact object (here a BH), we can notice broad and asymmetric reflection profiles in the spectrum due to distortion due to high gravitation. Proper modeling of the composite spectrum using different combinations of models can give good information about the physics that is going on there. In this regard, high signal-to-noise (s/n) data in the broadband range, e.g., Insight-HXMT (Zhang et al. 2020) can be very useful.
@@
Apart from the spectral features and variabilities, there are temporal variabilities also. Time variability, which is proportional to the length scale of the photon-emitting region, is as important as X-ray spectroscopy to understand the physical properties of accretion flow dynamics in the region near to the BH. During an outburst, we notice that there are very small time-scale variabilities in the light curves, especially in the high-energy ones. If we make a Fourier transformation of the light curve, we notice that fast variability is present in the power density spectrum or PDS (van der Klis 1989). In the PDS, which is a variation of $P(\nu)$ with $\nu$, the variability is mainly aperiodic, which is a direct consequence that the accretion flow onto the Bh is turbulent itself. The aperiodic variability, which is the result of the combination of the variation of the source flux and the variation in Poisson statistics, contains both broadband noise and narrow features. The broadband noise is spread over a large frequency range, whereas the narrow feature is a power peak in narrow frequency ranges. The narrow peak is known as the quasi-periodic oscillation (QPO), which is one of the most important properties of a BH. There is a break frequency ($\nu_{b}$) that also characterizes the PDS (van der Klis 1994). The noise is modelled by the power-law function, whereas the QPOs are modelled by one or multiple Lorentzian profiles. In BHs, we see low-frequency QPOs (LFQPOs) mainly due to their geometrical origin. Depending on some properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al. 2005). Although the clear origin of QPOs is still a matter of debate at present, several ideas were put forward as to why these features are being observed. Explanations like Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999), shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015) were put forward to explain the origin of LFQPOs in stellar-mass BHs. The shock oscillation model can give a simultaneous explanation of both the spectral and temporal properties, which makes it a very generally comprehensive model. High-resolution light curve data becomes very useful to study the temporal variability properties of BH. In this regard, Insight-HXMT is also very useful as well as studying the spectral properties.
@@
Apart from these properties and features of a BH outburst, there are a few other estimations that give a direct correspondence between the spectral and temporal properties. A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram or HID (Homan et al. 2001). Some model-fitted approaches are also important in this regard, e.g., accretion rate ratio or ARR, accretion rate ratio intensity diagram or ARRID (Mondal et al. 2014; Jana et al. 2016; Chatterjee et al. 2020). The RMS-intensity diagram or RID (Munoz-Darias et al. 2011), and hardness ration-intensity diagram or HRD (Belloni et al. 2005) are also very important estimates when it comes to correspondence between spectral and temporal properties. In different spectral states, all of these properties show different values, which are quite general for BH outbursts.
@@
The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023). The source was very bright as MAXI (Matsuoka et al. 2009) detected a flux of $7~{}Crab$ in the $2-20$ keV energy range. After that multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al. 2023; Miller-Jones et al. 2023; Negoro et al. 2023b; O’Connor et al. 2023; Wang & Bellm 2023; Williams-Baldwin et al. 2023). The source is reported to have a distance of $2.7\pm 0.3kpc$ (MataSanchez et al. 2024) and a mass of $10\pm 2~{}M_{\odot}$ (Svoboda et al. 2024). The revelation of the presence of hydrogen and helium emission lines in the optical spectrum lead to its classification as an LMXRB (Castro-Tirado et al. 2023). The X-ray spectrum and other properties point to its nature as a BHXRB (Liu et al. 2023; Sunyaev et al. 2023). This was supported strongly after the detection of type C QPOs (Palmer & Parsotan 2023; Draghis et al. 2023; Bollemeijer et al. 2023; Mereminskiy et al. 2023) and flat radio spectrum (Miller-Jones et al. 2023; Bright et al. 2023). There is X-ray monitoring of the source with many satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc. Using IXPE data on September 7, 2023, Veledina et al. (2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and PA of $32^{\circ}.2\pm 1^{\circ}.$3. There was a further polarization study of the source by Ingram et al. (2023). The presence of evolving LFQPOs at the high energy band of using AstroSat data was reported by Nandi et al. (2024). Using X-ray spectroscopy, there were reports about the spectral nature of the source. At the onset of the outburst, the spectral nature was consistent with a hard state with a photon index of power-law ($\Gamma$) $\sim 1.5-1.7$ (Liu et al. 2023). Then using the NICER observations, Bollemeijer et al. (2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days of the start of the outburst. The presence of LFQPOs was reported by Draghis et al. (2023). Using NICER data on August 29 2023 (MJD 60185), Debnath et al. (2023) reported the presence of a soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV at QPO frequency. With reflection spectroscopy, they found the inclination to be $\sim 87^{\circ}$.
@@
Here is how this paper unfolds. In §1, we give an introduction. In §2, we describe the observation, data selection, reduction, and analysis. In §3, we portray the results. In §4, we make the discussion about the possible physical origin of our result. And finally, in §5, we make a summary of our results and draw some conclusions.
@@ 2 Observation, Data Selection, Reduction, and Analysis
This source is being extensively monitored by various X-ray satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc. For our analysis, we use China’s first dedicated X-ray satellite Insight-High X-ray Modulation Telescope (HXMT; Zhang et al. 2020). In the subsequent subsections, we discuss the data selection, reduction, and analysis, respectively.
@@ 2.1 Data Selection
This source was regularly monitored by the Insight-HXMT satellite. In the data archive, observations are available on-demand. At the time of the work, we found that there was a total of 34 observation IDs were available, although 3 of them were not downloadable. Out of those usable 31 observation IDs, we took the first 10 observation IDs. However, unlike other X-ray satellite data, each observation ID in HXMT has several different exposures. Also, some of the observation IDs continue for almost 2 to 3 days. We found that for these first 10 observation IDs, each of them has a minimum of 7 to a maximum of 18 exposures. We checked the analysis of all the exposures for the first 3 observation IDs and found that no significant change was noticed. Thus we did not fit every exposure for each observation ID. The details about the selected data are given in Table 1.
Table 1: List of Data used
Obs. Id.[1] Start Date[2] End Date[2] Start MJD[3] Stop MJD[3] Exp. (s)[4]
@@
[1] Column 1 lists each observation ID.
[2] In columns 2 and 3, we give the start and end date of each observation ID.
[3] In columns 4 and 5, we give the start and end MJD of each observation ID.
[4] Column 6 gives the exposure time of each observation ID.

### v2

@@ Abstract
The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time that started in August 2023 and lasted for almost 6 months. We study the timing and spectral properties of this source using publicly available archival Insight-HXMT data for the first 10 observation IDs that last from MJD 60181 to 60198 with a total of 92 exposures for all three energy bands. We extracted the quasi-periodic oscillation properties by model fitting the power density spectrum and from those properties we designate that the QPOs are type-C in nature. We also conclude that the origin of the QPOs could be the shock instabilities in the transonic advective accretion flows around black holes. The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of Insight-HXMT in the broad energy band of $2-150$ keV. To achieve the best fit, spectral fitting required a combination of models e.g. interstellar absorption, power-law, multi-color disk-blackbody continuum, gaussian emission/absorption, and reflection by neutral material. From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was making a transition toward the softer states. The inner edge of the accretion disk moved inward in progressive days following the spectral nature. We found that the source has a high inclination. The average hydrogen column density estimated from the model fitting is $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
@@ 1 Introduction
Stellar-mass black holes (SBHs) are one of the end products of the death of massive stars. They reside in binary systems where a companion star supplies matter and during accretion they emit radiation. A significant amount of their radiation comes out in the form of X-rays, therefore, they are called black hole X-ray binaries (BHXRBs). Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs). The LMXRBs consist of type-A or later type of stars, whereas the HMXRBs consist of giant O or B-type of stars (White et al. 1995; Tetarenko et al. 2016). During the transfer of mass from the companion, matter accumulates at a distance from the BH due to insufficient viscosity to drive the matter inward, which is called the pile-up radius (Chakrabarti et al. 2019; Chatterjee et al. 2022). Due to the accumulation of more matter, the temperature will rise. The gradual increase in temperature develops instability at the pile-up radius, which further increases the viscosity. At some point in time when the accumulated matter gains enough viscosity to push forward, accretion starts. On the other hand, irradiation of the accumulated matter at the outer boundary by the central object can also trigger mass accretion when the temperature of the accumulated matter crosses a critical limit (King & Ritter 1998; Mondal 2020). Thus, these systems make ideal laboratories for studying the physics of accretion. The BHXRBs mostly stay in a quiet dormant state when their flux remains at a very low detection level. Episodically, the radiation level increases by a very large factor and they become easily detectable, which is known as an outburst (Remillard & McClintock 2006). The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence. Depending on the nature of the outburst, the sources are classified into two types: persistent and transient sources. Most of the BHXRBs are transient type of sources. They stay in the dormant state for most of their lifetimes. The luminosity becomes $\sim 10^{35}$ erg s-1 during episodic outbursting phase.
@@
During the onset of an outburst, the flux in the light curve changes noticeably. Based on its nature, the outbursts are divided into two types (Debnath et al. 2010): fast rise slow decay (FRSD) and slow rise slow decay (SRSD). However, there are also some other classified outburst types, e.g., according to Zhang et al. (2019), outbursts can also be classified into glitch, reflare, multipeak, and mini-outburst. There are generally four designated spectral states during a complete BH outburst (Remillard & McClintock 2006). They are known as the hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS). Generally, when an outburst starts, it starts in the HS. Then it slowly moves toward the SS through the HIMS and SIMS (Belloni et al. 2005, 2011, and references within). After it reaches its SS, the rising phase generally comes to an end and the outburst starts its decaying phase. In the decaying phase, it goes in the opposite direction and reaches the declining HS. When a BH outburst goes through all these states, it is known as a type-I outburst. When the soft state is absent, it is known as a ‘failed’ or type-II outburst.
The radiation spectrum of a BHXRB generally consists of two main components. One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973). The other is the hard power-law component, which can extend up to very high energies. The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which is intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al. 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al. 1999). When the outburst is in HS, there is a very small contribution from the thermal blackbody component and mostly the (non)thermal powerlaw component dominates. In SS, the thermal blackbody component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable. Other than these two components, in the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a closer to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Chakrabarti & Titarchuk 1995; Psaltis & Lamb 1997; Borozdin et al. 1999; Psaltis 2001). It was recently observed by Chatterjee et al. (2023) that BMC was present in the radiation spectrum of the black hole candidate (BHC) MAXI J0637-430.
In addition, reprocessing of the Comptonized photons from the Compton cloud by the accretion disk is known as reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010), which generates reflection spectra. As a result, the Fe $K\alpha$ emission line at $\sim 6.4$ keV and a reflection hump above 20 keV may originate. Therefore, the presence of complex features in the spectrum in BHXRBs makes them important to study to understand their origin. This can be achieved by fitting broadband data using multicomponent models. In this regard, high-resolution broadband data of Insight-HXMT (Zhang et al. 2020) can be very useful.
@@
Besides rich spectral features and variabilities, temporal properties are equally important to understanding the dynamics of the accreting gas around the BHs. During an outburst, it has been observed that the light curves show very small timescale variabilities, especially in the high-energy bands. The Fourier transformation of the light curve imprints such variabilities as both broadband noise and narrow features in the power density spectrum or PDS (van der Klis 1989). The broadband noise is spread over a large frequency range modeled by power-law function. The narrow feature is a power peak in narrow frequency ranges, known as the quasi-periodic oscillation (QPO), modeled by one or multiple Lorentzian profiles. In BHXRBs, the low-frequency QPOs (LFQPOs) are often observed mainly due to their geometrical origin. Depending on their properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), (%)RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al. 2005). Although their origin is still a topic of debate, several ideas were put forward e.g., Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999), shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015). The shock oscillation model can explain both the spectral and temporal properties simultaneously, which makes it a more comprehensive model.
@@
As both the properties are coming from the same system, and changes in spectral states can also dictate the types of QPOs, it is believed that both are interlinked. A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram, or HID (Homan et al. 2001). Some model-fitted approaches also showed that the changes in both properties are due to the interplay between different mass accretion rates or variation in viscosity, can be understood through accretion rate ratio intensity diagram or ARRID (Mondal et al. 2014; Jana et al. 2016; Chatterjee et al. 2020). The RMS-intensity diagram or RID (Munoz-Darias et al. 2011), and hardness ratio-intensity diagram or HRD (Belloni et al. 2005) can also make a link between spectral and temporal properties from pure observational ground.
@@
The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023). MAXI (Matsuoka et al. 2009) detected the source to be very bright with a flux of $7$ Crab in the $2-20$ keV energy range. Since then multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al. 2023; Miller-Jones et al. 2023; Negoro et al. 2023b; O’Connor et al. 2023; Wang & Bellm 2023; Williams-Baldwin et al. 2023). The source is reported to be located at a distance of $2.7\pm 0.3$ kpc (MataSanchez et al. 2024) and a mass of $10\pm 2~{}M_{\odot}$ (Svoboda et al. 2024). The revelation of the presence of hydrogen and helium emission lines in the optical spectrum led to its classification as an LMXRB (Castro-Tirado et al. 2023). The X-ray spectrum and other properties point to its nature as a BHXRB (Liu et al. 2023; Sunyaev et al. 2023). This was supported strongly after the detection of type C QPOs (Palmer & Parsotan 2023; Draghis et al. 2023; Bollemeijer et al. 2023; Mereminskiy et al. 2023) and a flat radio spectrum (Miller-Jones et al. 2023; Bright et al. 2023). There is X-ray monitoring of the source by NICER, NuSTAR, Insight-HXMT, IXPE, etc. Using IXPE data on September 7, 2023, Veledina et al. (2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and polarization angle (PA) of $32^{\circ}.2\pm 1^{\circ}.$3. There was a further polarization study of the source by Ingram et al. (2023).
At the starting of the outburst, the spectral nature was consistent with a hard state with a power-law photon index ($\Gamma$) $\sim 1.5-1.7$ (Liu et al. 2023). Later, using the NICER observations, Bollemeijer et al. (2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days from the starting of the outburst. The presence of LFQPOs was reported by Draghis et al. (2023) and the evolution of LFQPOs at the high energy band using AstroSat data was reported by Nandi et al. (2024). The source also showed soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV on August 29 2023 (MJD 60185) (Debnath et al. (2023) using NICER data). With reflection spectroscopy, authors found the disk inclination to be $\sim 87^{\circ}$. However, to date, there is no detailed study on the origin of the spectral and temporal properties of the source, which prompted us to analyze the broadband X-ray data from Insight-HXMT.
@@
The paper is organized as follows: in §2, we describe the observation, data selection, reduction, and analysis procedures. In §3, we portray the spectral and temporal properties and observed results. In §4, we discuss the possible physical origin of our result and connection between observed features. Finally, in §5, we summarize the results and draw conclusions.
@@ 2 Observation, Data Selection, Reduction, and Analysis
The source is being extensively monitored by various X-ray satellites, as discussed earlier. For this study, we use China’s first dedicated X-ray satellite Insight-HXMT (HXMT; Zhang et al. 2014, 2020) data. In the subsequent subsections, we discuss the data selection, reduction, and analysis, respectively.
@@ 2.1 Data Selection
The source was regularly monitored by the Insight-HXMT satellite. In the data archive, observations are available on-demand. We found that there are a total of 34 observation IDs available  [^1] Although 3 of them were not downloadable. Out of those available 31 observation IDs, we took the first 10 observation IDs. However, unlike other X-ray satellite data, each observation ID in HXMT has several exposures. Also, some of the observation IDs have continuous observations of about 2 to 3 days. We found that for these first 10 observation IDs, each of them has a minimum of 7 to a maximum of 18 exposures. Initially, we perform the analysis of all the exposures for the first 3 observation IDs and found no significant change in the parameters of spectral fits in the consecutive exposures. Therefore, we avoid fitting every exposure of each observation ID. The details of the selected data are given in Table 1.
Table 1: List of observations used for this work. Here, column 1 lists the observation IDs. Columns 2 and 3 denote the start and end date of observations, while columns 4 and 5 refer to corresponding MJDs. Column 6 gives the exposure time of each observation.
Obs. Id. Start Date End Date Start MJD Stop MJD Exp. (s)